OpenStudy (anonymous):

Calculate the mass of the products produced by the reaction of 7.5 moles of Na and 7.63 moles of H2O. 2Na + 2H2O ----- 2 NaOH + H2 someone can answer this i will give medals

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

can u help me here buddy i will give you medals and i will be your fan

OpenStudy (abhisar):

awww..that's sweet...let's see what we got here.

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

So from the equation we can see that 2 moles of Na reacts with 2 moles of H2O to produce two moles of NaOH and one mole of H2, this means that 1 mole of Na will react with 1 mole of H20 to produce 1 mole of NaOH and 0.5 moles of H2

OpenStudy (abhisar):

is that ok ?

OpenStudy (anonymous):

how can we can calculate it?

OpenStudy (abhisar):

So, 7.5 moles of Na will react with 7.5 moles of H2O to produce 7.5 moles of NaOH and 3.25 moles of H2

OpenStudy (abhisar):

Calculate what ?

OpenStudy (anonymous):

the mass?

OpenStudy (abhisar):

Ok, now for calculating the final mass of the products, we will do as following. 7.5 moles X Molar mass of NaOH + 3.25 moles X molar mass of H2

OpenStudy (abhisar):

= 7.5 X 40 + 3.25 X 2 =300+6.5=306.5 g

OpenStudy (anonymous):

how did you get the 3.25?

OpenStudy (abhisar):

Because, 1 mole of Na reacts with 1 mole of H2O to form 0.5 moles of H2. So 7.5 moles of Na will react with 7.5 moles of H2O to form 3.25 moles of H2

OpenStudy (abhisar):

See the equation 2Na + 2H2O ----- 2 NaOH + H2

OpenStudy (anonymous):

 how about the 2 NaOH + H2...what is the moles of that?

OpenStudy (anonymous):

i need to find the limiting and excess reactant

OpenStudy (anonymous):

i dont get it -.-

OpenStudy (abhisar):

Yes, you are going correct. Na is limmiting reactant here.

OpenStudy (abhisar):

Since, Na is the limmiting reagent, the reaction will be determined by the quantity of Na. So despite the fact that there is 7.63 moles of H2O is present only 7.5 moles of H2O will react with 7.5 moles of Na